If we create a 4 digit pin number using only the digits {0,1,2,3,4},

how many unique 4 digit pin numbers can we make?
Answer #1

Total number of 4 digit pins possible = 5 * 4 * 3 * 2 = 120   
Answer #2

Problem Statement:
We need to determine how many unique 4-digit PIN numbers can be created using only the digits {0, 1, 2, 3, 4}.


Key Observations:

  1. Digits Allowed: The digits available are 0, 1, 2, 3, 4. This gives us 5 possible choices for each digit in the PIN.


  2. PIN Length: The PIN is 4 digits long.

  3. Repetition: The problem does not restrict repetition, so digits can be repeated (e.g., 0000, 1111, etc., are allowed).


  4. Leading Zeros: The problem allows the PIN to start with 0 (e.g., 0123), so it is a 4-digit sequence rather than a strict 4-digit number where the first digit cannot be 0.


Calculation:

  • For each of the 4 positions in the PIN, there are 5 possible choices (0, 1, 2, 3, 4).

  • Since the choices are independent, the total number of unique PINs is calculated by multiplying the number of choices for each digit:

    Total PINs=5×5×5×5=54Total PINs=625

Answer:
The number of unique 4-digit PIN numbers that can be created is 625


answered by: anonymous